This Question: 30 pts 7 of 7 (1 complete) This Test: 200 pts possible 0 Marias Foods produces frozen meals that it sells forThis Question: 30 pts 7 of 7 (1 complete) This Test: 200 pts possible Marias Foods produces frozen meals that it sells for $This Question: 30 pts 7 of 7 (1 complete) This Test: 200 pts possible o Marias Foods produces frozen meals that it sells for
Answer #1

1.

Product Cost
Absorption Costing Variable Costing
Variable Manufacturing Cost $            4.00 $           4.00
Fixed Manufacturing Overhead $            0.60
Total Product Cost $            4.60 $           4.00

2a.

Income Statement (Absorption)
Sales Revenue $          8,550 =950*9
Cost of Goods Sold $          4,370 =950*4.6
Gross Profit $          4,180
Selling and Administrative costs $          1,250 =950*1+300
Net Operating Income $          2,930

2b.

Income Statement (Variable)
Sales Revenue $          8,550 =950*9
Variable Costs $          4,750 =950*5
Contribution Margin $          3,800
Fixed Costs $             990 =690+300
Net Operating Income $          2,810

3.
Absorption Costing Income is higher

If you have any query, kindly comment with your query and please mark thumbs up.